OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2) A. y = 3x – 5 B. y = 3x – 13 C. y = 3x – 1 D. y = 3x + 2

OpenStudy (anonymous):

the line is parallel to the line y = 3x +2 so the slope is 3 and substitution (5,2) and slope of 3 into a basic quadratic equation y=kx+b so 5=3*2+b b=-1 so y=3x-1